Gamma–normal distribution (from http://www.math.wm.edu/˜ leemis/chart/UDR/UDR.html ) The details associated with gamma–normal distribution are given on page 112 of Forbes, Evans, Hastings, and Peacock (2011), Statistical Distributions, Fourth Edition, John Wiley and Sons. The gamma–normal distribution appears to be a bivariate, rather than a univariate distribution.